# Controllable Video Generation: A Survey

### v1

@@ 4.4.3 Camera-Guided Generation
Camera control signals can be provided as explicit sequences of camera poses (extrinsics: rotation $\mathbf{R}_{t}$ and translation $\mathbf{T}_{t}$, and intrinsics $\mathbf{K}_{t}$ if they vary), relative transformations between frames, or higher-level textual descriptions of camera movements (e.g., “dolly zoom out from the character”). Motion Prompting [5] explores textual descriptions for camera and object motion.
Recent advancements heavily leverage diffusion models conditioned on camera parameters. CameraCtrl [98] and its successor CameraCtrl II [244] provide explicit control over camera trajectories. Direct-a-Video [220], also mentioned for trajectory control, supports camera path guidance. ViewCrafter [256] and Cavia [255] focus on generating novel views based on camera inputs.
@@
CamCo [265] and Trajectory Attention [263] explore camera control in specific contexts. Uni3C [243] explores the joint control of human motion and camera trajectory. The camera parameters are typically used to define projection matrices that transform latent scene representations or guide the sampling of features to render the scene from the specified viewpoint. EgoSim [254] focuses on egocentric camera simulation. Approaches like I2VControl [210] , CamI2V [262], and RealCam-I2V [261] specialize in image-to-video generation with camera control. More advanced systems like 3DTrajMaster [246], MotionFlow [247], and CineMaster [253] aim for comprehensive 3D-aware camera and motion control.
@@
Challenges include maintaining 3D scene consistency and avoiding disocclusion artifacts or distorted geometry when the camera undergoes significant movement or viewpoint changes. Generating high-fidelity novel views that are consistent with previous frames is difficult. Ensuring smooth and natural camera transitions, free from jitter or abrupt changes, is crucial for cinematic quality. Providing intuitive user interfaces for specifying complex 3D camera paths and parameters remains an open area. Boosting Camera Motion [250], MotionMaster [251], OmniCam [268], and ReCamMaster [252] are among works aiming to improve the quality and control of camera movements. Techniques like VD3D [389, 216], Aether [6], and GenDoP [257] explore complex 3D scene understanding and camera path generation. The latest work Voyager [267] introduces a world caching scheme and smooth video sampling, and Follow-Your-Creation [266] utilizes the temporal pack inference strategy to keep the long-term consistency, respectively.
@@ 4.4.4 Motion-Guided Generation
The work generally utilizes an explicit motion signal as the control anchor, with an optional text prompt defining scene semantics. Motion control signals vary in format: structured pose sequences like MoTrans [280] and MotionPrompting [5], hand mask sequences like InterDyn [270], or optical flow fields like OnlyFlow [281] provide dense or sparse motion descriptors. Some works directly use reference videos as motion priors, including DreamMotion [286], DreamRelation [140], Customize-A-Video [287], Follow-Your-Motion [292], DualReal [168], and VideoMage [169]. Other works encode abstract fields, such as velocity maps [275] or relational cues. Even in motion-guided settings, text prompts are retained to define the appearance, layout, or objects involved in the scene. Furthermore, a growing subset of works supports zero-shot or training-free transfer, including MotionClone [282] and VMC [289], by leveraging alignment in latent spaces or score distillation techniques.
@@
To realize controllable generation, a range of architectural designs and mechanisms have been introduced, often combining motion-aware modules with large-scale diffusion backbones. UniAnimate DiT [269] and DreamActor-M1 [271] inject motion tokens into DiT transformers for coherent spatiotemporal control. InterDyn [270] performs targeted gesture synthesis via attention-guided injection of hand masks. DreamMotion [286] distills score maps from reference videos to enforce temporal behavior. DreamRelation [140] models trajectory consistency using relation fields between appearance regions. Methods like MotionAgent [275] and MotionPrompting [5] formulate motion as fields or anchor prompts for fine control, while SMA [279] and Diffusion as Shader [278] propose spectral and 3D-aware conditioning mechanisms. Meanwhile, MotionMatcher [274], Motionbooth [152], and MotionDirector [291] align motion tokens in latent space, and Customize-A-Video [287] learns a one-shot controller from minimal supervision.
@@
Despite these advances, motion-guided generation still faces significant challenges. Motion signals can be noisy, ambiguous, or misaligned with the model’s internal representation, often causing motion artifacts or flicker. Recent work MotionPro [293] proposes a region-wise trajectory and motion mask to alleviate misinterpretation and achieve better motion control.
@@
Furthermore, generalization to unseen motions or subjects remains limited, particularly in one-shot or zero-shot settings (e.g., MotionClone [282], VMC [289]). Evaluation also lags behind: although SST-EM [277] introduces metrics for semantic, spatial, and temporal fidelity, widely accepted benchmarks and interpretability tools are still under development.
@@ 4.5 Audio Control
Audio control in video generation refers to the ability to synthesize videos from audio signals such as speech, music, or general sound. This form of control enables models to generate temporally aligned and semantically coherent videos based on audio cues. A prominent subdomain is voice control, which focuses on generating talking face videos or personalized portrait animations conditioned on speech input. Another direction is sound control, which utilizes broader forms of audio (e.g. music or ambient sound) to guide the generation of more general and diverse video content beyond facial animation.
@@ 4.5.1 Voice-Guided Generation
Voice-guided video generation aims to use both audio and image as conditions to create realistic videos, especially for the talking portrait video generation tasks, which is considered one of the most crucial trends in the applications of the human-computer interaction field. The input embeddings of the sound condition sequences $S_{t}$ firstly are extracted commonly by the pretrained encoder like wav2vec [390]. In order to acquire richer semantic information from the input audio, several works [310],[311],[316] have concatenated features from different encoder layers for the $f$-th frame $S^{f}$ and injected them into the diffusion processes by the cross-attention layers.
@@
The core issue in talking portrait video generation tasks is capturing and estimating the actual face expressions from the various voice inputs, such as the movements of lips, head poses, and eyebrows, which influence the quality of the video significantly. For some earlier GAN-based methods, Wav2Lip [320] adapts a powerful lip-sync discriminator to generate an accurate, realistic lip motion, while Audio-driven Talking Face Video Generation [321] aims to address the problem of unnatural head movements for talking face generation application. OneShotA2V [322] proposes a framework capable of developing a talking face from an unseen image and an audio clip. Recently, VividTalk [317] and AniPortrait [315] propose a two-stage architecture based on 3D mesh as the intermediate representation. EDTalk [314] has disentangled the motion space into three distinct latent spaces and designs an audio-to-motion module to predict different expressions. InstructAvatar [312] accomplishes the fine-grained control based on the natural language instruction. EMO [316] designs the audio feature extractor, face locator, and speed layers, combining with the ReferenceNet [71] and motion frame module, which ensures the performance of generated videos. EchoMimic [4] and EchoMimicV2 [93] also use the ReferenceNet to extract the features from the input image, generating a better portrait video. Aiming to provide a precise lip control, DreamTalk [309], SayAnything [300], MuseTalk [297] and CAFE-TALK [298] propose optimized techniques to solve this problem. ANGIE [323] utilize VQ-Motion Extractor and Co-Speech GPT module to achieve co-speech gesture video generation.
@@
Different from other conditions like images, layouts, and sketches, the strength of sound is weaker, hindering the control effectiveness during the inference process. Therefore, addressing the conflict between various conditions is another challenge for voice-guided video generation. V-Express [311] proposes a progressive training strategy and conditional dropout operations, which is benefitial for those weak control conditions. MOFA-Video [211] designs a MOFA-Adapter to animate a reference image into a video under various control conditions. To balance the control strength between different conditions, MegActor-$\Sigma$  [307] proposes a mixed-modal DiT, ACTalker [295] introduced a mask-drop strategy, while MotionCraft [308] and OmniHuman-1 [302] optimize the training strategy of the multimodal generation framework.
@@
Besides those challenges, there are still some issues such as limitations in style variation generation on efficiency. For the former one, SVP [305] models the intrinsic style of the source video to alleviate this problem. For the latter one, FLOAT [304], VASA-1 [306], ChatAnyone [296] and CyberHost [303] improve the real-time performance for the video-chat application. Finally, as for the consistency in long-term video generation, MCDM [301] designs two clip-motion-prior modules and a memory-efficient attention mechanism. In contrast, TexTalker [299] focuses on the dynamic texture for the high-fidelity talking heads generation.
@@ 4.5.2 Sound-Guided Generation
In sound-guided video generation, sound inputs range from specific sounds like animal calls and music to general audio cues. Various audio feature extraction methods, such as Mel frequency cepstral coefficients (MFCCs), Mel spectrograms, BEATs, and CLAP, are used to capture texture, pitch, rhythm, and emotional tones. For instance, AV-Link [324] and ASVA [325] synchronize visuals with audio by extracting rhythmic, spectral, and emotional properties using BEATs, CLAP, and MFCCs or Mel-Spectrograms. TA2V [326] further integrates text, using BEATs to guide video generation, showcasing how audio dynamics shape visual output.
Early sound-guided video generation methods, such as TräumerAI [332], Sound2Video [329], and Sound2Sight [330], rely on GANs to generate videos from audio, but struggle with audio-video synchronization and fine control over motion. Later, CCVS [331] improves temporal consistency by integrating optical flow and contextual information, though it still requires additional control signals. Recent advancements in sound-guided video generation, such as AV-Link [324] and ASVA [325], have shifted toward diffusion models for more accurate synchronization, using temporally-aligned activations and audio features to ensure better alignment. TA2V [326] further enhances this by incorporating both text and audio inputs, conditioning video motion and emotion. MotionCraft [308] introduces a dual-branch architecture for rhythmic and semantic alignment, while DAA2V [327] refines the process with text-to-video models for improved temporal and semantic alignment. MagicInfinite [328] further advances these methods by offering precise audio-lip synchronization and fine-grained control over facial animations using a two-stage learning approach.
Challenges in sound-guided video generation include maintaining synchronization between audio and video, especially with diverse sound types. Solutions like AV-Link[324] and ASVA[325] use temporal attention mechanisms to address this. TA2V[326] tackles the complexity of coordinating both text and audio, while DAA2V[327] provides insights on maintaining synchronization across different audio types.

### v2

@@ 5.4.3 Camera-Guided Generation
Camera control signals can be provided as explicit sequences of camera poses (extrinsics: rotation $\mathbf{R}_{t}$ and translation $\mathbf{T}_{t}$, and intrinsics $\mathbf{K}_{t}$ if they vary), relative transformations between frames, or higher-level textual descriptions of camera movements (e.g., “dolly zoom out from the character”). Motion Prompting [geng2024motionprompting] explores textual descriptions for camera and object motion. Recent advancements heavily leverage diffusion models conditioned on camera parameters. CameraCtrl [he2024cameractrl] and its successor CameraCtrl II [he2025cameractrl2] provide explicit control over camera trajectories. Direct-a-Video [yang2024directavideo], also mentioned for trajectory control, supports camera path guidance. ViewCrafter [yu2024viewcrafter] focuses on generating novel views based on camera inputs. CamCo [xu2024camco] and Trajectory Attention [xiao2024trajectoryattention] explore camera control in specific contexts. Uni3C [cao2025uni3cunifyingprecisely3denhanced] explores the joint control of human motion and camera trajectory. The camera parameters are typically used to define projection matrices that transform latent scene representations or guide the sampling of features to render the scene from the specified viewpoint. EgoSim [yu2025egosim] focuses on egocentric camera simulation. Approaches like I2VControl [feng2024i2vcontrol] , CamI2V [zheng2024cami2v], and RealCam-I2V [li2025realcami2v] specialize in image-to-video generation with camera control. More advanced systems like 3DTrajMaster [fu2024trajmaster], MotionFlow [zhang2025motionflow], and CineMaster [wang2025cinemaster] aim for comprehensive 3D-aware camera and motion control.
@@
Challenges include maintaining 3D scene consistency and avoiding disocclusion artifacts or distorted geometry when the camera undergoes significant movement or viewpoint changes. Generating high-fidelity novel views that are consistent with previous frames is difficult. Ensuring smooth and natural camera transitions, free from jitter or abrupt changes, is crucial for cinematic quality. Providing intuitive user interfaces for specifying complex 3D camera paths and parameters remains an open area. ReCamMaster [bai2025recammaster] aims to improve the quality and control of camera movements. Techniques like VD3D [bahmani2025vd3d], Aether [zhu2025aether], and GenDoP [zhang2025gendop] explore complex 3D scene understanding and camera path generation. The latest work Voyager [huang2025voyagerlongrangeworldconsistentvideo] introduces a world caching scheme and smooth video sampling, and Follow-Your-Creation [ma2025followcreation] utilizes the temporal pack inference strategy to keep the long-term consistency, respectively.
Table 6: Quantitative comparisons of camera-guided video generation on T2V setting across RealEstate10K and DL3DV. Frame Consistency and Motion Strength are denoted as FC and MS. In addition. The rotation and translation error are denoted as RE and TE.
Method RealEstate10K DL3DV
FVD $\downarrow$ CLIPSIM $\uparrow$ FC $\uparrow$ MS $\uparrow$ RE $\downarrow$ TE $\downarrow$ FVD $\downarrow$ CLIPSIM $\uparrow$ FC $\uparrow$ MS $\uparrow$ RE $\downarrow$ TE $\downarrow$
MotionCtrl [wang2024motionctrl] 506.9 0.2744 0.9471 0.9734 2.90 1.06 746.7 0.2033 0.9448 0.9462 1.17 3.23
CameraCtrl [he2024cameractrl] 426.8 0.2734 0.9526 0.9750 2.68 0.98 675.5 0.2027 0.9518 0.9610 1.01 4.00
AC3D [bahmani2025ac3d] 415.6 0.3044 0.9496 0.9811 2.67 0.38 452.8 0.2078 0.9529 0.9869 1.06 2.11
DualCamCtrl [zhang2025dualcamctrl] 408.1 0.3154 0.9540 0.9918 1.23 0.25 427.4 0.2090 0.9544 0.9807 0.83 1.53
@@ 5.4.4 Motion-Guided Generation
The work generally utilizes an explicit motion signal as the control anchor, with an optional text prompt defining scene semantics. Motion control signals vary in format: structured pose sequences like MoTrans [li2024motrans] and MotionPrompting [geng2024motionprompting], hand mask sequences like InterDyn [akkerman2025interdyn], or optical flow fields like OnlyFlow [koroglu2024onlyflow] provide dense or sparse motion descriptors. Some works directly use reference videos as motion priors, including DreamMotion [jeong2024dreammotion], DreamRelation [wei2025dreamrelation], Customize-A-Video [ren2024customizeavideo], DualReal [wang2025dualrealadaptivejointtraining], and VideoMage [Huang_2025_CVPR]. Other works encode abstract fields, such as velocity maps [liao2025motionagent] or relational cues. Even in motion-guided settings, text prompts are retained to define the appearance, layout, or objects involved in the scene. Furthermore, a growing subset of works supports zero-shot or training-free transfer, including MotionClone [ling2024motionclonetrainingfreemotioncloning] and VMC [jeong2023vmc], by leveraging alignment in latent spaces or score distillation techniques.
@@
To realize controllable generation, a range of architectural designs and mechanisms have been introduced, often combining motion-aware modules with large-scale diffusion backbones. DreamActor-M1 [luo2025dreamactor] injects motion tokens into DiT transformers for coherent spatiotemporal control. InterDyn [akkerman2025interdyn] performs targeted gesture synthesis via attention-guided injection of hand masks. DreamMotion [jeong2024dreammotion] distills score maps from reference videos to enforce temporal behavior. DreamRelation [wei2025dreamrelation] models trajectory consistency using relation fields between appearance regions. Methods like MotionAgent [liao2025motionagent] and MotionPrompting [geng2024motionprompting] formulate motion as fields or anchor prompts for fine control, while SMA [pondaven2024spectralmotion] and Diffusion as Shader [gu2025diffusionasshader] propose spectral and 3D-aware conditioning mechanisms. Meanwhile, MotionMatcher [wu2025motionmatcher], Motionbooth [wu2024motionbooth], and MotionDirector [zhao2023motiondirector] align motion tokens in latent space, and Customize-A-Video [ren2024customizeavideo] learns a one-shot controller from minimal supervision.
@@
Despite these advances, motion-guided generation still faces significant challenges. Motion signals can be noisy, ambiguous, or misaligned with the model’s internal representation, often causing motion artifacts or flicker. Recent work MotionPro [Zhang_2025_CVPR] proposes a region-wise trajectory and motion mask to alleviate misinterpretation and achieve better motion control.
@@
Furthermore, generalization to unseen motions or subjects remains limited, particularly in one-shot or zero-shot settings (e.g., MotionClone [ling2024motionclonetrainingfreemotioncloning], VMC [jeong2023vmc]). Evaluation also lags behind: although SST-EM [biyyala2025sstem] introduces metrics for semantic, spatial, and temporal fidelity, widely accepted benchmarks and interpretability tools are under development.
5.4.5 Summary
Temporal Control activates the dynamic layer of the framework, employing motion modules (e.g., Trajectory Samplers) and camera control modules to translate static concepts into coherent narratives. This modality is central to satisfying the user needs for automation and augmentation, as it replaces manual keyframing with semantic guidance like flow fields or drag-based interactions (e.g., DragAnything [wu2024draganything]), thereby streamlining the creation of complex scenes.
Regarding the architecture, while UNet models (e.g., Motion-I2V [10.1145/3641519.3657497]) excel at short-term pixel-level flow dynamics, the complexity of Video-to-4D generation is necessitating a move towards DiT (e.g., DreamActor-M1 [luo2025dreamactor]) and AR backbones. These architectures are favored not just for size, but for their ability to model long-term causal dependencies and 3D geometric consistency (e.g., Voyager [huang2025voyagerlongrangeworldconsistentvideo]), which are critical when coordinating sophisticated camera paths with object interactions over extended durations.
In the future, the trend of temporal control is moving towards integrated control where temporal signals (Trajectories/Flow) need to coexist with spatial controls (Structure/ID). Current general control models often struggle to merge these distinct signals seamlessly, highlighting a user demand for better interactive editing tools (e.g., drag-based interfaces like DragAnything [wu2024draganything]) that provide real-time preview of motion paths before generation. Future frameworks will likely prioritize universal control mechanisms to unify these diverse motion inputs into a single interface.
[…]
Method Video Quality & Coherence Audio Fidelity & Quality Audio-Visual Synchronization
[…]
MM-Diffusion [ruan2023mm] 0.32 0.43 0.13 0.99 - 5.37 4.07 4.27 5.89 - - - - - 0.12
JavisDiT [liu2025javisdit] 0.34 0.53 0.38 0.99 0.38 5.46 2.24 3.19 4.54 1.00 0.14 0.89 11.62 1.13 0.18
UniVerse-1 [wang2025universe] 0.52 0.67 0.24 0.99 0.89 5.52 2.13 3.63 4.84 0.24 0.07 0.97 10.71 1.10 0.12
Ovi [low2025ovi] 0.57 0.65 0.34 0.99 0.90 6.19 2.13 4.44 5.84 0.49 0.12 4.04 9.62 1.14 0.18
Harmony [hu2025harmony] 0.59 0.65 0.36 0.99 0.91 6.39 2.05 4.73 5.67 0.15 0.12 5.61 7.53 0.92 0.19
@@ 5.5 Audio Control
Audio control in video generation refers to the ability to synthesize videos from audio signals such as speech, music, or general sound. This form of control enables models to generate temporally aligned and semantically coherent videos based on audio cues. A subdomain is voice control, which focuses on generating talking face videos or personalized portrait animations conditioned on speech input. Another direction is sound control, which utilizes broader forms of audio (e.g., music or ambient sound) to guide the generation with more general content beyond facial animation.
@@ 5.5.1 Voice-Guided Generation
Voice-guided video generation aims to use both audio and image as conditions to create realistic videos, especially for the talking portrait video generation tasks, which is considered one of the most crucial trends in the applications of the human-computer interaction field. The input embeddings of the sound condition sequences $S_{t}$ firstly are extracted commonly by the pretrained encoder like wav2vec [schneider2019wav2vecunsupervisedpretrainingspeech]. In order to acquire richer semantic information from the input audio, several works [xu2024hallohierarchicalaudiodrivenvisual, wang2024vexpressconditionaldropoutprogressive, tian2024emoemoteportraitalive] have concatenated features from different encoder layers for the $f$-th frame $S^{f}$ and injected them into the diffusion processes by the cross-attention layers.
@@
The core issue in talking portrait video generation tasks is capturing and estimating the actual face expressions from the various voice inputs, such as the movements of lips, head poses, and eyebrows, which influence the quality of the video significantly. Recently,  [sun2023vividtalk][wei2024aniportraitaudiodrivensynthesisphotorealistic] propose a two-stage architecture based on 3D mesh as the intermediate representation. EDTalk [tan2024edtalkefficientdisentanglementemotional] has disentangled the motion space into three distinct latent spaces and designs an audio-to-motion module to predict different expressions. EMO [tian2024emoemoteportraitalive] designs the audio feature extractor, face locator, and speed layers, combining with the ReferenceNet [hu2024animateanyoneconsistentcontrollable] and motion frame module, which ensures the performance of generated videos. EchoMimic [chen2024echomimiclifelikeaudiodrivenportrait] and EchoMimicV2 [meng2024echomimicv2strikingsimplifiedsemibody] also use the ReferenceNet to extract the features from the input image, generating a better portrait video.
@@
Different from other conditions like images, layouts, and sketches, the strength of sound is weaker, hindering the control effectiveness during the inference process. Therefore, addressing the conflict between various conditions is another challenge for voice-guided video generation. V-Express [wang2024vexpressconditionaldropoutprogressive] proposes a progressive training strategy and conditional dropout operations, which are beneficial for those weak control conditions. MOFA-Video [niu2024mofavideo] designs a MOFA-Adapter to animate a reference image into a video under various control conditions. To balance the control strength between different conditions, MegActor-$\Sigma$  [yang2024megactorsigmaunlockingflexiblemixedmodal] proposes a mixed-modal DiT, ACTalker [hong2025audiovisualcontrolledvideodiffusion] introduced a mask-drop strategy, while MotionCraft [bian2025motioncraft] and OmniHuman-1 [lin2025omnihuman1rethinkingscalinguponestage] optimize the training strategy of the multimodal generation framework.
@@
Besides those challenges, there are still some issues, such as limitations in style variation generation on efficiency. For the former one, SVP [tan2024svpstyleenhancedvividportrait] models the intrinsic style of the source video to alleviate this problem. For the latter one,  [ki2024floatgenerativemotionlatent, xu2024vasa1lifelikeaudiodriventalking, lin2025cyberhost] improve the real-time performance for the video-chat application. Finally, as for the consistency in long-term video generation, MCDM [shen2025longtermtalkingfacegenerationmotionprior] designs two clip-motion-prior modules and a memory-efficient attention mechanism. In contrast, TexTalker [li2025highfidelity3dtalkingavatar] focuses on the dynamic texture for the high-fidelity talking heads generation.
@@ 5.5.2 Sound-Guided Generation
In sound-guided video generation, sound inputs range from specific sounds like animal calls and music to general audio cues. Various audio feature extraction methods, such as Mel frequency cepstral coefficients (MFCCs), Mel spectrograms, BEATs, and CLAP, are used to capture texture, pitch, rhythm, and emotional tones. For instance, AV-Link [avlink] and ASVA [Zhang2024AudioSynchronizedVA] synchronize visuals with audio by extracting rhythmic, spectral, and emotional properties using BEATs, CLAP, and MFCCs or Mel-Spectrograms. TA2V [zhao2024ta2v] further integrates text, using BEATs to guide video generation, showcasing how audio dynamics shape visual output.
Recent advancements in sound-guided video generation, such as AV-Link [avlink] and ASVA [Zhang2024AudioSynchronizedVA], have shifted toward diffusion models for more accurate synchronization, using temporally-aligned activations and audio features to ensure better alignment. TA2V [zhao2024ta2v] further enhances this by incorporating both text and audio inputs, conditioning video motion and emotion. MotionCraft [bian2025motioncraft] introduces a dual-branch architecture for rhythmic and semantic alignment, while DAA2V [yariv2024diverse] refines the process with text-to-video models for improved temporal and semantic alignment.
Challenges in sound-guided video generation include maintaining synchronization between audio and video, especially with diverse sound types. Solutions like AV-Link[avlink] and ASVA[Zhang2024AudioSynchronizedVA] use temporal attention mechanisms to address this. TA2V[zhao2024ta2v] tackles the complexity of coordinating both text and audio, while DAA2V[yariv2024diverse] provides insights on maintaining synchronization across different audio types.
5.5.3 Summary
In this domain, audio functions as a temporal-dynamic anchor, guiding the rhythmic progression of visual content. A primary challenge is fine-grained temporal alignment to prevent synchronization drift. This issue has driven a shift from UNet-based pipelines (e.g., EchoMimic [chen2024echomimiclifelikeaudiodrivenportrait]) to unified joint generation. Spearheading this evolution, recent works like MegActor-$\Sigma$ [yang2024megactorsigmaunlockingflexiblemixedmodal] leverage Diffusion Transformers (DiT) to ensure superior long-range coherence.
Driven by users’ demand for high-fidelity interaction, current research tackles a critical control imbalance, where dominant visual features frequently suppress subtle audio cues. It necessitates balanced training strategies, such as the conditional dropout in V-Express [wang2024vexpressconditionaldropoutprogressive]. However, the separation of structural voice and semantic sound control remains a limitation, suggesting a future move toward holistic audio-visual physics to unify micro and macro dynamics.